Fabina borrows ₹ 12,500 at 12% per annum for 3 years at simple interest and Radha borrows the same amount for the same time period at 10% per annum, compounded annually. Who pays more interest and by how much? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Interest paid by Fabina =P x R x T
100
=12500 x 3 x 12
100
= 4500
Amount paid by Radha at the end of 3 years =A =P (1 +R/100 )^n
=₹ [12500(1+ 10/100 )^3]
=₹ [12500 × (110/100)^3]
=₹ 16637.50
C.I. = A − P = ₹ 16637.50 − ₹12500 =₹4,137.50
The interest paid by Fabina is₹4,500 and by Radha is₹4,137.50.
∴ Fabina pays more interest.
₹ 4500 − ₹4137.50 =₹362.50
Hence, Fabina will have to pay₹362.50 more

​Thanks